# Arena: A Patch-of-Interest ViT Inference Acceleration System for Edge-Assisted Video Analytics

### v1

@@ 7. Conclusion
In this paper, we find visual foundation models like ViT also have a dedicated acceleration mechanism for video analytics. Then, we design Arena, a novel edge-assisted, real-time inference acceleration system tailored for ViT-based models. This system aims to accelerate video analytics by reducing the number of patches offloaded from cameras to the server and alleviating the computation overhead of the model while still maintaining high accuracy. Our results show that Arena can accelerate the inference up to 1.58×1.58$\times$1.58 × and 1.82×1.82$\times$1.82 × on average by only using 54% and 34% bandwidth while maintaining a high inference accuracy.

### v2

@@ VII Conclusion
In this paper, we find visual foundation models like ViT also have a dedicated acceleration mechanism for video analytics. Then, we design Arena, a novel edge-assisted, real-time inference acceleration system tailored for ViT-based models. Arena periodically operates in keyframe inference and non-keyframe inference. The PPS algorithm innovatively guides the filtering of redundant information in video frames during keyframe inference. MFR is effectively used to reconstruct unstructured feature maps in non-keyframes. The AKIS algorithm switches inference strategies based on the content of the video. This system aims to accelerate video analytics by reducing the number of patches offloaded from cameras to the server and alleviating the computation overhead of the model while still maintaining high accuracy. We deploy Arena on the prototype that uses an NVIDIA Jetson as a camera device and an Ubuntu desktop as an edge server running real video analytics workloads. Our results show that Arena can accelerate the inference up to 1.58×1.58$\times$1.58 × and 1.82×1.82$\times$1.82 × on average by only using 47% and 31% bandwidth while maintaining a high inference accuracy. In the future, we will extend Arena to a broader range of downstream tasks.